OpenStudy (anonymous):

write the equation of the line that has slope2 and that passes through the point (1,3).

OpenStudy (anonymous):

slope=m=2 general form of a straight line is: y=mx+c we hav m=2,x=1 and y=3 we get c=1 therefore equation is y=2x+1.

OpenStudy (anonymous):

or u can start from y-3=2(x-1) Point slope form

OpenStudy (anonymous):

2x-y+1=0